OpenStudy (karolynxx):

The point-slope form of the equation of the line that passes through (–4, –3) and (12, 1) is y – 1 = 1/4 (x – 12). What is the standard form of the equation for this line? a. x – 4y = 8 b. x – 4y = 2 c. 4x – y = 8 d. 4x – y = 2

OpenStudy (dayakar):

multiply both sides with 4

OpenStudy (dayakar):

can u do it

OpenStudy (dayakar):

4y-4 = 4*1/4(x-12) 4y-4 = x-12 add -4y both sides

OpenStudy (karolynxx):

makes no sense to me

OpenStudy (dayakar):

-4y+4y-4 = x-4y-12 -4 = x-4y-12 add 12 both sides 12-4 =x-4y-12+12 8 = x-4y x-4y=8 answer is a

OpenStudy (xmissalycatx):

Here's an example of how to get to standard form! :)

OpenStudy (karolynxx):

Thank you!